OpenStudy (anonymous):

Jackson wants to purchase flowers to plant on Mother's day for his mom. If each plant costs $1.25 and he has $10 to spend, write an equation for the amount of money, y, Jackson will have left over after purchasing x amount of plants

OpenStudy (anonymous):

Maybe this would be easier to see if you had easier numbers. If you had $5, and each plant costs $2, how would you find out how much money, y, you would have left over after buying x = 2 plants?

OpenStudy (anonymous):

Actually, for you particular problem, it's even easier... If each plant costs $1.25, then # plants cost per plant total cost left over from $10 1 $1.25 1.25 10 - 1.25 = 8.75 2 $1.25 2.50 10 - 2.50 = 7.50 3 $1.25 3.75 10 - 3.75 = 6.25 etc... So the amount left over from $10 is ten dollars minus (1.25 times the number of plants)

OpenStudy (anonymous):

since the amount left over is called "y" and the number of plants purchased is called "x" Another way to write the sentence I wrote above... "the amount left over from $10 is ten dollars minus (1.25 times the number of plants)" is like this: y = 10 - 1.25x amount left = 10 minus (1.25 times # of plants) over